OpenStudy (anonymous):

The functions f(x) = 32x + 8 and g(x) are described below: f(x) = 32x + 8 g(x) = 32x - 9 The graph of g(x) is obtained by shifting down the graph of f(x) by ______ units?

OpenStudy (anonymous):

Notice that to change f(x) into g(x) all we have to do to f(x) is subtract 17 f(x) - 17 = 32x + 8 - 17 = 32x - 9 = g(x)

OpenStudy (anonymous):

When you subtract a number from a function, it shifts it down by that many units.

OpenStudy (anonymous):

Thank you so much! I am checking my son's algebra test and I have forgotten everything. He got that one right!